OpenStudy (anonymous):

Find the polynomial function with roots 11 and 2i.

OpenStudy (anonymous):

all i got was x-11)(x-2i)(x+2i)=0 (x-11)(x^2-4i^2) (x-11)(x^2+4) x^3-11x^2+4x-44=0

OpenStudy (anonymous):

can some on please tell me if im right